# Attack Anything: Blind DNNs via Universal Background Adversarial Attack

### v1

@@ Abstract
Index Terms:
Deep neural networks, adversarial perturbation, background adversarial attack, convergence analysis, attack anything.
I Introduction
The remarkable advancements of deep learning have revolutionized various domains of artificial intelligence (AI), enabling significant achievements in computer vision, natural language processing, and other complex tasks [1]. However, these achievements have also unveiled a critical vulnerability of deep neural networks (DNNs) to adversarial perturbations [2, 3, 4, 5, 6, 7, 8, 9]. Numerous studies [10, 11, 12, 13, 14, 15, 16, 17] have demonstrated the alarming ease with which state-of-the-art (SOTA) models can be manipulated through carefully crafted perturbations, raising great concerns about DNNs’ reliability and security.
@@ I Introduction
Figure 1: Comparison of adversarial perturbations in diverse forms. (a) conducts digital attacks with imperceptible perturbations entirely covering the images [2]. (b)-(h) perform physical attacks by corrupting targeted objects with physical perturbations in various forms [3, 18, 19, 20, 21, 23, 11]. (i) is our adversarial attack with background perturbation preserving the integrity of the targeted objects. Please zoom in for the details of the background attack.
Existing studies [24, 25, 26, 27, 28] have primarily centered on adversarial attacks that corrupt targeted objects (physical attack) or images (digital attack) as shown in Fig. 1 (a)-(h). These attacks are designed to be ”visually” camouflaged for DNNs, a strategy that is intuitively plausible and comprehensible given that humans can also be deceived by visually camouflaged objects. However, an interesting divergence arises when considering the impact of background variations on the targeted objects. While such variations do not significantly affect human recognition, DNNs exhibit a high degree of sensitivity to these changes, as exemplified by the banana and donut in Fig. 1 (i). This discrepancy underscores a fundamental difference in the role of background features in human and machine vision. Historically, adversarial attacks have overlooked the potential of exploiting background features, resulting in an incomplete understanding of their role in adversarial contexts. Moreover, the prevailing focus on a specific object (physical attack) or whole image (digital attack) manipulation may not sufficiently address the need for generalizing adversarial attacks. These limitations impede progress in exploring the adversarial robustness of DNNs.
@@
In this paper, we redirect the attention toward background adversarial attacks that are executed smoothly across digital and physical domains, transferring well across various objects, models, and tasks. By manipulating the background environment without directly interfering with objects, we introduce a novel approach to adversarial attacks, i.e., we propose an innovative framework, capitalizing on the untapped potential of background features to deceive DNNs, as shown in Fig. 2. Methodologically, we formulate the background adversarial attack as an iterative optimization problem, analogous to the process of DNN learning, and provide a theoretical demonstration of its convergence under certain moderate but sufficient conditions. To enhance the attack transferability and efficacy, we introduce a novel ensemble strategy tailored to the unique attributes of adversarial perturbations, effectively strengthening their capability in various scenarios. Additionally, we propose a sophisticated smooth constraint that ensures the harmonious integration of perturbations. To validate the efficacy and robustness of the proposed method, we undertake an extensive series of experiments. These experiments span across both the digital and physical realms, white-box and black-box conditions, involving diverse objects, models, and tasks. The experimental results underscore the formidable effectiveness of the introduced background adversarial attack framework, revealing its potential to disrupt a wide range of AI applications in real-world scenarios. The implications of our findings extend beyond the realm of adversarial attacks, prompting a profound reevaluation of the principles that underpin DNNs.
In summary, our contributions are as follows:
@@
The remainder of this paper is organized as follows. Section II briefly reviews adversarial attacks and convergence analysis concerning DNNs. Section III details the proposed universal background adversarial attack. Section IV presents the experimental results and analyses. Section V discusses the implications of the findings. Section VI concludes the paper.
II Backgrounds
In this section, we give the backgrounds of adversarial attacks according to different attack domains (II-A and II-B) and convergence analysis concerning DNNs (II-C).
II-A Digital Attack
@@
II-B Physical Attack
Physical adversarial attacks, in contrast, extend the concept of adversarial attacks into the physical realm. The primary motivation behind physical attacks is to craft physical modifications, causing the deep learning models to be misinterpreted. Numerous AI systems have fallen under physical attacks, such as face recognition [40, 41], autonomous driving [42, 18], remote sensing [43, 44], etc. Researchers have demonstrated that by applying adversarially designed stickers [23, 45], patterns [46, 47], makeup [48, 49], light [50, 51], 3D mesh [11], etc., to an object, DNNs-based AI systems can misidentify the object as something entirely different. However, the aforementioned physical attacks share a commonality in that they all need to corrupt the targets of interest in varying forms. Some studies [52, 53, 54] have endeavored to manipulate the backgrounds of targeted objects for adversarial purposes, causing slight sway in the model’s predictions, yet often devoid of comprehensive empirical substantiation. Additionally, a fraction of these effects might stem from data augmentations beyond the model’s training regimen. Research [55] proposes a contextual background attack (CBA) against aerial detection by crafting adversarial patterns tailored for aircraft, which achieves comparable performance while lacking comprehensive reflection of the potential of background attack.
II-C Convergence Analysis
Convergence analysis is a critical aspect of studying DNNs. It involves understanding how the iterative learning process of a DNN progresses and whether it will eventually reach a point where the model’s parameters no longer change significantly, indicating that the model has learned the underlying patterns in the training data. Yang et al. [56] first explore the convergence of training DNNs with stochastic momentum methods, in particular for non-convex optimization, which fills the gap between practice and theory by developing a basic convergence analysis of two stochastic momentum methods. Work [57] provides a fine-grained convergence analysis for a general class of adaptive gradient methods including AMSGrad [58], RMSProp [59] and AdaGrad [60]. The authors of [58] fix the convergence issue of Adam-type algorithms by endowing them with long-term memory of past gradients. In paper [61], the researchers develop an analysis framework with sufficient conditions, which guarantee the convergence of the Adam-type methods for non-convex stochastic optimization
@@ II-C Convergence Analysis
III Methodology
In this section, we first formulate the problem of background adversarial attack in III-A and give a detailed illustration of the proposed paradigm of attack anything in III-B. Then we describe the ensemble strategy in III-C and objective loss in III-D for attacking anything, respectively. Finally, we conduct a convergence analysis of the devised background attack in III-E.
@@ III Methodology
III-A Problem Formulation
Previous studies have predominantly focused on carrying out adversarial attacks by directly corrupting targeted objects or images. These attacks aim to ”visually” blind DNNs, which is intuitively feasible and understandable since humans can also be deceived by visually camouflaged objects. However, an interesting divergence arises when considering the impact of background variations. While such variations hardly affect human recognition, DNNs exhibit a high degree of sensitivity to these changes. This discrepancy underscores a fundamental difference in the role of background features in the visual perception of humans and machines. Historically, adversarial attacks have overlooked the potential of exploiting background features, resulting in an incomplete understanding of their role in adversarial contexts. In contrast, this paper redirects the focus toward background adversarial attacks that can easily blind DNNs even without causing any disruptions to the targeted objects themselves.
@@ III-A Problem Formulation
where $D(\cdot)$ is parameterized with $\boldsymbol{\theta}$, $\varnothing$ means recognition results are no-objects or background. To achieve the aforementioned attack purpose, we construct the objective loss of the background adversarial attack as $L(D(\boldsymbol{x}^{*},\boldsymbol{\theta}),\boldsymbol{x}^{*})$, which is concretely explained in Sec. III-D. We mathematically formulate this attack as an optimization problem similar to training a DNN as follows:
@@
7:     $[\boldsymbol{x}^{1}_{i},\boldsymbol{y}^{1}_{i},\boldsymbol{x}^{2}_{i},\boldsymbol{y}^{2}_{i},\boldsymbol{s}^{conf}_{i},\boldsymbol{p}^{cls}_{i}]\leftarrow D(\boldsymbol{x}^{*}_{i})$;
@@ III-B Attack Anything
$\mathbf{g}_{t}&=\nabla\sum\limits_{b=1}^{B_{t}}L(D(\boldsymbol{x},\boldsymbol{M}_{objs},\boldsymbol{M}_{bg},\boldsymbol{P}^{(t)},\boldsymbol{\theta}),\boldsymbol{P}^{(t)})\\ &=\nabla f_{t}({\boldsymbol{P}^{(t)}}),\end{aligned}\end{split}$ (7)
@@ III-C Ensemble Strategy
III-D Objective Loss
III-D1 Adversarial Loss
@@ III-E Convergence Analysis
$\leq$ $\frac{\max_{i}\left(G_{i}\right)}{\sum_{t=1}^{T}\alpha_{t}}\cdot\left(\left(\frac{L}{2}\frac{\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\sum_{i=1}^{d}G_{i}^{2}/c^{2}\right.\right.$
@@
$\left.+\left(L\cdot 2\frac{\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\left(\max_{i}G_{i}\right)^{2}\frac{\alpha_{1}}{\left(1-\beta_{1}\right)c}\right.\right.$
$\left.\left.+\frac{\beta_{1}}{1-\beta_{1}}\left(\max_{i}G_{i}\right)\left(\max_{i}G_{i}\right)\right.\right.$
@@
(a) Bounding box loss ($L_{box}$)
(b) Ojectiveness loss ($L_{obj}$)
(c) Total variation loss ($L_{abtv}$)
(d) Total loss ($L$)
As a consequence, our formulated background adversarial attack is mathematically convergent with mild sufficient conditions, which is also demonstrated with experimental results as shown in Fig. 7. The loss functions are detailed in Sec. III-D. Through the above convergence analysis, we made a positive step toward understanding the theoretical behavior of the proposed background attack methods.
IV Experiments
In this section, we present the experimental settings in IV-A. Then, we demonstrate the effectiveness of attacking anything in both digital and physical domains in IV-B and IV-C, respectively. Furthermore, we compare the proposed background adversarial attack with SOTA physical attacks in IV-D. Next, we showcase the effectiveness of attacking anything across different objects, models, and tasks in IV-E. Finally, we conduct an ablation study to verify the effectiveness of the proposed ensemble strategy and novel smoothness loss in IV-F.
@@ IV Experiments
TABLE I: Experimental results of digital background attack on the validation set of COCO in the metric of mAP, where white-box attacks are highlighted in bold and the rest are black-box attacks. The redder the cell, the worse the detection performance. The bluer the cell, the better the detection performance. Clean and Random Noise mean experiments on clean images and images with random noise, respectively. The 19 detectors of the first row and the first column are for detection and perturbation optimization, respectively.
IV-A Experimental Settings
IV-A1 Models
@@ IV-A4 Implementations
IV-B Digital Background Attacks
We perform digital background attacks with numerous mainstream object detection methods. Specifically, we use the validation set of COCO to verify digital attack efficacy by replacing the objects’ backgrounds with the elaborated adversarial perturbations. We report the quantitative experimental results in Table I, and the metric is mAP0.5. In addition, we also visualize the quantitative experimental results from the perspective of detection and attack in terms of ASR in Fig. 8. It is demonstrated that:
@@ IV-B Digital Background Attacks
TABLE II: Experimental results of physical background attack in the metric of DR, where white-box attacks are highlighted in bold and the rest are black-box attacks. The redder the cell, the higher the attack efficacy. The bluer the cell, the lower the attack efficacy. Clean and Random Noise mean experiments on clean images and images with random noise, respectively. The 19 detectors of the first row and the first column are for detection and perturbation optimization, respectively.
@@
IV-C Physical Background Attacks
@@
TABLE III: Quantitative attack comparison of car detection in physically-based simulation in the metric of DR, where the best results are highlighted in bold. The redder the cell, the higher the attack efficacy. The bluer the cell, the lower the attack efficacy. TH and RN mean the threshold of confidence score and random noise, respectively. ”fg”, ”bg”, and ”bf” represent the perturbation on foreground, background, and both, respectively.
@@
TABLE IV: Quantitative attack comparison of person detection in physically-based simulation in the metric of DR, where the best results are highlighted in bold. The redder the cell, the higher the attack efficacy. The bluer the cell, the lower the attack efficacy. TH and RN mean the threshold of confidence score and random noise, respectively. ”fg”, ”bg”, and ”bf” represent the perturbation on foreground, background, and both, respectively.
(a) Attack perspective
(b) Detection perspective
We report the quantitative experimental results in Table II, and the metric is DR. In addition, we also visualize the quantitative experimental results from the perspective of detection and attack in terms of ASR in Fig. 10. It is concluded that:
IV-D Physical Attack Comparison
The quantitative experimental results of car and person detection are shown in Table III and Table IV, respectively. In addition, we also display the qualitative experimental results of car and person detection in the Appendix. The confidence score lines of the correct detection are shown in Fig. 11 and Fig. 12 to further illustrate the attack performance. It is observed that:
IV-E Attack Anything
IV-E1 Across Different Models
As shown in Table I and II, the method generalizes well across various models in the white box and black box conditions for most cases. However, some perturbations generated by detectors with similar structures may transfer well between each other, while it is hard to generalize to other models as shown in Table II. The devised ensemble attack may properly resolve the above issues. The experimental results are shown in Table. VI. It is observed that the attack transferability is significantly improved by our designed ensemble strategy.
IV-E2 Across Different Objects
We conduct physical attacks on YOLOv5 with different objects, such as a bottle, person, cup, car, and several kinds of fruits. The quantitative experimental results are exhibited in Table V. We can observe that the proposed attack anything framework generalizes well between various objects with DR decreasing to 0 for most cases. The qualitative experimental results as shown in Fig. 2 (d).
(a) Clean
(b) Random
(c) DTA
(d) FCA
(e) ACTIVE
(f) AA-fg
(g) AA-bg
(h) AA-bf
(a) Clean
(b) Random
(c) DTA
(d) FCA
(e) ACTIVE
(f) AA-fg
(g) AA-bg
(h) AA-bf
@@ IV-E2 Across Different Objects
Threshold 0.15 0.25 0.35 0.45 0.55
Bottle 0.000 0.000 0.000 0.000 0.000
Person 0.132 0.016 0.000 0.000 0.000
Apple 0.070 0.020 0.020 0.000 0.000
Banana 0.000 0.000 0.000 0.000 0.000
Orange 0.000 0.000 0.000 0.000 0.000
Cup 0.000 0.000 0.000 0.000 0.000
Car 0.168 0.045 0.018 0.000 0.000
TABLE V: The physical attack performance across different objects in the metric of DR with different thresholds of confidence score.
@@
TABLE VI: Ablation study on ensemble strategy (”Swin Transformer+YOLOv5m” and ”YOLOv5m+Mask R-CNN”) in physical attack settings in the metric of DR, where white-box attacks are highlighted in bold and the rest are black-box attacks. The redder the cell, the higher the attack efficacy. The bluer the cell, the lower the attack efficacy. The 19 detectors of the first row and the first column are for detection and perturbation optimization, respectively.
@@
Not smooth 0.762 0.573 0.420 0.322 0.185
@@
TABLE VII: Ablation study on smoothness setting with various thresholds of confidence score in the metric of DR.
IV-E3 Across Different Tasks

### v3

@@ Abstract
keywords:
DNNs , Universal , Background , Adversarial Attack
[…]
1 Introduction
The remarkable advancements of deep learning have revolutionized various domains of artificial intelligence (AI), enabling significant achievements in computer vision, natural language processing, and other complex tasks. However, these achievements have also unveiled a critical vulnerability of deep neural networks (DNNs) to adversarial perturbations [1, 2, 3, 4, 5, 6, 7]. Numerous studies [8, 9, 10, 11, 12, 13] have demonstrated the alarming ease with which state-of-the-art (SOTA) models can be manipulated through carefully crafted perturbations, raising great concerns about DNNs’ reliability and security.
@@ 1 Introduction
Figure 1: Comparison of adversarial perturbations in diverse forms. (a) conducts digital attacks with imperceptible perturbations entirely covering the images [1]. (b)-(h) perform physical attacks by corrupting targeted objects with physical perturbations in various forms [14, 15, 16, 17, 18, 20, 8]. (i) is our adversarial attack with background perturbation preserving the integrity of the targeted objects.
Existing studies [21, 22, 23, 24, 25] have primarily centered on adversarial attacks that corrupt targeted objects (physical attack) or images (digital attack) as shown in Fig. 1 (a)-(h). These attacks are designed to be “visually” camouflaged for DNNs, a strategy that is intuitively plausible and comprehensible given that humans can also be deceived by visually camouflaged objects. However, an interesting divergence arises when considering the impact of background variations on the targeted objects. While such variations do not significantly affect human recognition, DNNs exhibit a high degree of sensitivity to these changes, as exemplified by the banana and donut in Fig. 1 (i). This discrepancy underscores a fundamental difference in the role of background features in human and machine vision. Historically, adversarial attacks have overlooked the potential of exploiting background features, resulting in an incomplete understanding of their role in adversarial contexts. Moreover, the prevailing focus on a specific object (physical attack) or whole image (digital attack) manipulation may not sufficiently address the need for generalizing adversarial attacks. These limitations impede progress in exploring the adversarial robustness of DNNs.
@@
In this paper, we redirect the attention toward background adversarial attacks that are executed smoothly across digital and physical domains, transferring well across various objects, models, and tasks. By manipulating the background environment without directly interfering with objects, we introduce a novel approach to adversarial attacks, i.e., we propose an innovative framework, capitalizing on the untapped potential of background features to deceive DNNs, as shown in Fig. 2. Methodologically, we formulate the background adversarial attack as an iterative optimization problem, analogous to the process of DNN learning, and provide a theoretical demonstration of its convergence under certain moderate but sufficient conditions. To enhance the attack transferability and efficacy, we introduce a novel ensemble strategy tailored to the unique attributes of adversarial perturbations, effectively strengthening their capability in various scenarios. Additionally, we propose a sophisticated smooth constraint that ensures the harmonious integration of perturbations. To validate the efficacy and robustness of the proposed method, we undertake an extensive series of experiments. These experiments span across both the digital and physical realms, white-box and black-box conditions, involving diverse objects, models, and tasks. The experimental results underscore the formidable effectiveness of the introduced background adversarial attack framework, revealing its potential to disrupt a wide range of AI applications in real-world scenarios. The implications of our findings extend beyond the realm of adversarial attacks, prompting a profound reevaluation of the principles that underpin DNNs. In summary, our contributions are as follows:
@@
The remainder of this paper is organized as follows. Section 2 briefly reviews adversarial attacks and convergence analysis concerning DNNs. Section 3 details the proposed universal background adversarial attack. Section 4 presents the experimental results and analyses. Section 5 discusses the implications of the findings. Section 6 concludes the paper.
2 Backgrounds
In this section, we give the backgrounds of adversarial attacks according to different attack domains (2.1 and 2.2) and convergence analysis concerning DNNs (2.3).
2.1 Digital Attack
@@
2.2 Physical Attack
Physical adversarial attacks, in contrast, extend the concept of adversarial attacks into the physical realm. The primary motivation behind physical attacks is to craft physical modifications, causing the deep learning models to be misinterpreted. Numerous AI systems have fallen under physical attacks, such as face recognition [37, 38], autonomous driving [39, 15], remote sensing [40, 41], and so on. Researchers have demonstrated that by applying adversarially designed stickers [20, 42], patterns [43, 44], makeup [45, 46], light [47, 48], 3D mesh [8], etc., to an object, DNNs-based AI systems can misidentify the object as something entirely different. However, the aforementioned physical attacks share a commonality in that they all need to corrupt the targets of interest in varying forms. Some studies [49, 50, 51] have endeavored to manipulate the backgrounds of targeted objects for adversarial purposes, causing slight sway in the model’s predictions, yet often devoid of comprehensive empirical substantiation. Additionally, a fraction of these effects might stem from data augmentations beyond the model’s training regimen. Research [52] and [53] propose to perform background attack on aerial detection, which achieves comparable performance while lacking theoretical demonstration and comprehensive reflection of the potential of background attack.
2.3 Convergence Analysis
Convergence analysis is a critical aspect of studying DNNs. It involves understanding how the iterative learning process of a DNN progresses and whether it will eventually reach a point where the model’s parameters no longer change significantly, indicating that the model has learned the underlying patterns in the training data. Yang et al. [54] first explore the convergence of training DNNs with stochastic momentum methods, in particular for non-convex optimization, which fills the gap between practice and theory by developing a basic convergence analysis of two stochastic momentum methods. Work [55] provides a fine-grained convergence analysis for a general class of adaptive gradient methods including AMSGrad [56], RMSProp [57] and AdaGrad [58]. The authors of [56] fix the convergence issue of Adam-type algorithms by endowing them with long-term memory of past gradients. In paper [59], the researchers develop an analysis framework with sufficient conditions, which guarantee the convergence of the Adam-type methods for non-convex stochastic optimization.
@@ 2.3 Convergence Analysis
3 Methodology
In this section, we first formulate the problem of background adversarial attack in 3.1 and give a detailed illustration of the proposed paradigm of attack anything in 3.2. Then we describe the ensemble strategy in 3.3 and objective loss in 3.4 for attacking anything, respectively. Finally, we conduct a convergence analysis of the devised background attack in 3.5.
@@ 3 Methodology
3.1 Problem Formulation
Previous studies have predominantly focused on carrying out adversarial attacks by directly corrupting targeted objects or images. These attacks aim to “visually” blind DNNs, which is intuitively feasible and understandable since humans can also be deceived by visually camouflaged objects. However, an interesting divergence arises when considering the impact of background variations. While such variations hardly affect human recognition, DNNs exhibit a high degree of sensitivity to these changes. This discrepancy underscores a fundamental difference in the role of background features in the visual perception of humans and machines. Historically, adversarial attacks have overlooked the potential of exploiting background features, resulting in an incomplete understanding of their role in adversarial contexts. In contrast, this paper redirects the focus toward background adversarial attacks that can easily blind DNNs even without causing any disruptions to the targeted objects themselves.
@@ 3.1 Problem Formulation
where $D(\cdot)$ is parameterized with $\boldsymbol{\theta}$, $\varnothing$ means recognition results are no-objects or background. To achieve the aforementioned attack purpose, we construct the objective loss of the background adversarial attack as $L(D(\boldsymbol{x}^{*},\boldsymbol{\theta}),\boldsymbol{x}^{*})$, which is concretely explained in Sec. 3.4. We mathematically formulate this attack as an optimization problem similar to training a DNN as follows:
@@
7:  $\boldsymbol{x}^{1}_{i},\boldsymbol{y}^{1}_{i},\boldsymbol{x}^{2}_{i},\boldsymbol{y}^{2}_{i},\boldsymbol{s}^{conf}_{i},\boldsymbol{p}^{cls}_{i}]\leftarrow D(\boldsymbol{x}^{*}_{i})$;
@@ 3.2 Attack Anything
$\begin{split}\begin{aligned} \mathbf{g}_{t}&=\nabla\sum\limits_{b=1}^{B_{t}}L(D(\boldsymbol{x},\boldsymbol{M}_{objs},\boldsymbol{M}_{bg},\boldsymbol{P}^{(t)},\boldsymbol{\theta}),\boldsymbol{P}^{(t)})\\ &=\nabla f_{t}({\boldsymbol{P}^{(t)}}),\end{aligned}\end{split}$ (7)
@@ 3.3 Ensemble Strategy
3.4 Objective Loss
3.4.1 Adversarial Loss
@@ 3.5 Convergence Analysis
[…]
@@
[…]
@@
Table 1: Experimental results of digital background attack on the validation set of COCO in the metric of mAP, where white-box attacks are highlighted in bold and the rest are black-box attacks. The redder the cell, the worse the detection performance. The bluer the cell, the better the detection performance. Clean and Random Noise mean experiments on clean images and images with random noise, respectively. The 19 detectors of the first row and the first column are for detection and perturbation optimization, respectively.
(a) Bounding box loss ($L_{box}$)
(b) Ojectiveness loss ($L_{obj}$)
(c) Total variation loss ($L_{abtv}$)
(d) Total loss ($L$)
As a consequence, our formulated background adversarial attack is mathematically convergent with mild sufficient conditions, which is also demonstrated with experimental results as shown in Fig. 7. The loss functions are detailed in Sec. 3.4. Through the above convergence analysis, we made a positive step toward understanding the theoretical behavior of the proposed background attack methods.
4 Experiments
In this section, we present the experimental settings in 4.1. Then, we demonstrate the effectiveness of attacking anything in both digital and physical domains in 4.2 and 4.3, respectively. Furthermore, we compare the proposed background adversarial attack with SOTA physical attacks in 4.4. Next, we showcase the effectiveness of attacking anything across different objects, models, and tasks in 4.5. Finally, we conduct an ablation study to verify the effectiveness of the proposed ensemble strategy and novel smoothness loss in 4.6.
4.1 Experimental Settings
4.1.1 Models
@@ 4.1.4 Implementations
4.2 Digital Background Attacks
We perform digital background attacks with numerous mainstream object detection methods. Specifically, we use the validation set of COCO to verify digital attack efficacy by replacing the objects’ backgrounds with the elaborated adversarial perturbations. We report the quantitative experimental results in Table 1, and the metric is mAP0.5. In addition, we also visualize the quantitative experimental results from the perspective of detection and attack in terms of ASR in Fig. 8. It is demonstrated that:
@@ 4.2 Digital Background Attacks
Table 2: Experimental results of physical background attack in the metric of DR, where white-box attacks are highlighted in bold and the rest are black-box attacks. The redder the cell, the higher the attack efficacy. The bluer the cell, the lower the attack efficacy. Clean and Random Noise mean experiments on clean images and images with random noise, respectively. The 19 detectors of the first row and the first column are for detection and perturbation optimization, respectively.
@@
4.3 Physical Background Attacks
@@
(a) Attack perspective
(b) Detection perspective
We report the quantitative experimental results in Table 2, and the metric is DR. In addition, we also visualize the quantitative experimental results from the perspective of detection and attack in terms of ASR in Fig. 10. It is concluded that:
(a) Clean
(b) Random
(c) DTA
(d) FCA
(e) ACTIVE
(f) AA-fg
(g) AA-bg
(h) AA-bf
(a) Clean
(b) Random
(c) DTA
(d) FCA
(e) ACTIVE
(f) AA-fg
(g) AA-bg
(h) AA-bf
@@
Table 3: Quantitative attack comparison of car detection in physically-based simulation in the metric of DR, where the best results are highlighted in bold. The redder the cell, the higher the attack efficacy. The bluer the cell, the lower the attack efficacy. TH and RN mean the threshold of confidence score and random noise, respectively. “fg”, “bg”, and “bf” represent the perturbation on foreground, background, and both, respectively.
@@
Table 4: Quantitative attack comparison of person detection in physically-based simulation in the metric of DR, where the best results are highlighted in bold. The redder the cell, the higher the attack efficacy. The bluer the cell, the lower the attack efficacy. TH and RN mean the threshold of confidence score and random noise, respectively. “fg”, “bg”, and “bf” represent the perturbation on foreground, background, and both, respectively.
@@
Table 5: Ablation study on ensemble strategy (“Swin Transformer+YOLOv5m” and “YOLOv5m+Mask R-CNN”) in physical attack settings in the metric of DR, where white-box attacks are highlighted in bold and the rest are black-box attacks. The redder the cell, the higher the attack efficacy. The bluer the cell, the lower the attack efficacy. The 19 detectors of the first row and the first column are for detection and perturbation optimization, respectively.
@@
Bottle 0.000 0.000 0.000 0.000 0.000
Person 0.132 0.016 0.000 0.000 0.000
Apple 0.070 0.020 0.020 0.000 0.000
Banana 0.000 0.000 0.000 0.000 0.000
Orange 0.000 0.000 0.000 0.000 0.000
Cup 0.000 0.000 0.000 0.000 0.000
Car 0.168 0.045 0.018 0.000 0.000
Table 6: The physical attack performance across different objects in the metric of DR with different thresholds of confidence score.
Threshold 0.15 0.25 0.35 0.45 0.55
Unsmooth 0.762 0.573 0.420 0.322 0.185
@@
Table 7: Ablation study on smoothness setting with various thresholds of confidence score in the metric of DR.
4.4 Physical Attack Comparison
The quantitative experimental results of car and person detection are shown in Table 3 and Table 4, respectively. In addition, we also display the qualitative experimental results of car and person detection in the Appendix. The confidence score lines of the correct detection are shown in Fig. 11 and Fig. 12 to further illustrate the attack performance. It is observed that:
4.5 Attack Anything
4.5.1 Across Different Models
As shown in Table 1 and 2, the method generalizes well across various models in the white box and black box conditions for most cases. However, some perturbations generated by detectors with similar structures may transfer well between each other, while it is hard to generalize to other models as shown in Table 2. The devised ensemble attack may properly resolve the above issues. The experimental results are shown in Table. 5. It is observed that the attack transferability is significantly improved by our designed ensemble strategy.
4.5.2 Across Different Objects
We conduct physical attacks on YOLOv5 with different objects, such as a bottle, person, cup, car, and several kinds of fruits. The quantitative experimental results are exhibited in Table 6. We can observe that the proposed attack anything framework generalizes well between various objects with DR decreasing to 0 for most cases. The qualitative experimental results as shown in Fig. 2 (d).
4.5.3 Across Different Tasks